Find the slope and the equation of the tangent of the curve at the given point. Show your solution.



1) y = 2x² -x + 2 (-1, 5)



2) y = x² + x + 1 at (-2, 3)

1
Expert's answer
2022-03-16T06:49:33-0400

equation of the tangent of the curve at the point a can be found the following way

"f(x)-f(a)=f'(a)(x-a)" , f'(a) is the slope of this tangent line

1) y = 2x² -x + 2 (-1,5)

"f(-1)=5"

"f'(x)=4x-1\implies f'(-1)=-5" - slope

So, the sought equation is "f(x)-5=-5(x+1)\implies f(x)=-5x"


2) y = x² + x + 1 at (-2, 3)

"f(-2)=3"

"f'(x)=2x+1\implies f'(-2)=-3" - slope

So, the sought equation is "f(x)-3=-3(x+2)\implies f(x)=-3x-3"